Tom and Jerry had a basketball match. At the end of the match, the total score was 120 points. If tom scored 21 more points than half of the points that jerry scored, what was the final score of each team?
Answer by jorel1380(3719)

Let j be the points Jerry scored. Then Tom would be j/2 +21. So:
j+j/2 +21=120
2j+j+42=240
3j=198
j=66
j/2 +21=54
☺☺☺☺